# Unlearning via Sparse Representations

### v1

@@ 5.1 Experimental Setup
Unlearning
We aim to make the problem of unlearning as challenging as possible in order to fairly evaluate the proposed methods. Therefore, on each dataset we select the class that is best learned by the respective models with the Discrete Key Value Bottleneck trained previously, to be the forget class (see Appendix A for further details). Table 1 further shows the accuracy of the base models on the train and test splits of the thus defined retain and forget class dataset splits.
Objective & Metrics
We report our results on the test data of retain classes and forget class, i.e. $\mathcal{D}_{test}^{retain}$ and $\mathcal{D}_{test}^{forget}$. Further, in our experiments, we aim to achieve complete unlearning - achieving minimal accuracy on the forget set while maintaining the performance on the retain set. [^1] All of our experiments are performed on a RTX8000 GPU with 48GB memory. While achieving complete unlearning may not always be desirable, such as in the case of Membership Inference Attacks (MIAs) the proposed methods can be easily extended to defend against MIAs (We refer to Appendix D for further discussion on MIAs and the proposed methods).
5.2 Zero-Shot Unlearning via Discrete Key-Value Bottleneck
We will now discuss the results of unlearning via activations and examples, i.e. the two approaches proposed in Section 4 on all three benchmark datasets. All experiments are performed across 5 random seeds and mean values are reported.
Unlearning via Activations.
As discussed in Section 4, unlearning via activations requires us to set the hyperparameter $N_{a}$, reflecting the top-$N_{a}$ most frequently activated key-value pairs which will be masked out after inference on the forget set. We therefore start by analyzing its role over a wide range of values of $N_{a}$ to probe its choice and effect including $N_{a}=0$ being the limit without any unlearning. Figures 2(c) - 2(e) summarize the unlearning and effect of $N_{a}$ on the retain vs forget test set. In the case of CIFAR-10, the initial accuracies on the retain and forget test set are 92.61% and 96.50% respectively. As $N_{a}$ increases, the forget class test accuracy decreases, slowly for small $N_{a}$ and rapidly for larger $N_{a}$. The model reaches random accuracy (i.e. 10% for CIFAR-10) on the forget class test data at $N_{a}=150000$. At this point the retain set test accuracy is 92.97%. The model unlearns the forget class completely between $N_{a}=170,000$ (0.4%) and $N_{a}=180,000$ (0%). At this point, the retain set test accuracy is 92.94%, which is almost identical to the initial accuracy. Further increasing $N_{a}$ up to $N_{a}=200,000$, i.e. about 20% of all key-value pairs, leads to an additional increase in retain set test accuracy to 93%. As can be seen from the equivalent analysis on the CIFAR-100 models in Figure 2(d) as well as the Lacuna-100 models in Figure 2(e), the same trend of maintaining the initial retain accuracy while minimizing the forget accuracy up to a minimum holds across all three datasets validating its meaningful unlearning capability.
(c) CIFAR-10
(d) CIFAR-100
(e) LACUNA-100
Figure 2: Performance on the retain set test data vs. Performance on the forget set test data on (a) CIFAR-10 (b) CIFAR-100 (c) LACUNA-100 in the case of Unlearning via Activations as the value of $N_{a}$ is increased. The relative performance on the retain set test data as compared to the original models increases after unlearning in the case of CIFAR-10 and drops by 0.2% and 0.17% for CIFAR-100 and LACUNA-100.
Unlearning via Examples.
For the second method—unlearning via examples—$N_{e}$ examples are sampled randomly from the training data of the forget class, and subsequently used for unlearning by the mechanism described in Section 4. Similar to before, we aim to assess the effect on the choice of $N_{e}$ over a wide range for each dataset, including $N_{e}=0$ being the limit without any unlearning. Figures 3(a) - 3(c) summarize the unlearning and effect of $N_{a}$ on the retain vs. forget test set. We again begin by focusing on the results with CIFAR-10 (Figure 3(a)). Here, the forget set $\mathcal{D}_{train}^{forget}$ contains 5000 examples. We start off with retain set and forget set test accuracies of 92.61% and 96.50% respectively. Similar to the previous approach – unlearning via activations – the test accuracy on the forget set decreases with increasing $N_{e}$. The accuracy on the retain test set, on the other hand, increases monotonically, although only slightly overall. The model achieves random accuracy on the forget class around $N_{e}=2500$. The accuracy on retain set test data is at just under 93% at this transition. Finally, the accuracy on the forget set drops to 0% (i.e. complete unlearning) between $N_{e}=3000$ and $N_{e}=3400$ with a retain set test accuracy of just above 93% at $N_{e}=3400$. Further increasing $N_{e}$ does not affect the retain set test accuracy notably. An equivalent analysis on the CIFAR-100 models in Figure 3(b) and the Lacuna-100 models in Figure 3(c) exhibits a very similar trend to CIFAR-10 with the only difference that the initial retain accuracy is roughly maintained instead of further increasing. Successful minimization of the forget accuracy up to a minimum is achieved across all three datasets, validating it as another option for unlearning using discrete key-value bottlenecks.
(a) CIFAR-10
(b) CIFAR-100
(c) LACUNA-100
Figure 3: Performance on the retain set test data vs. Performance on the forget set test data on (a) CIFAR-10 (b) CIFAR-100 (c) LACUNA-100 in the case of Unlearning via Examples as the value of $N_{e}$ is increased. Similar to Unlearning via Activations, the relative performance on the retain set test data as compared to the original models increases after unlearning in the case of CIFAR-10 and drops by 0.36% and 0.09% for CIFAR-100 and LACUNA-100.
Summary.
Both methods, Unlearning via Activations and Unlearning via Examples, successfully demonstrated unlearning of the forget class while having a negligible effect on the models’ performance on the retain set. Importantly, this is achieved without any form of training, retraining, or fine-tuning as is usually required by other methods. The retain set test accuracy remains more or less constant for all three datasets except for a few minor fluctuations. This is a result of the fact that due to localized and context-dependent sparse updates during the initial training of the model, discrete key-representations corresponding to different classes in the dataset are well separated from each other, an important prerequisite discussed in (Träuble et al., 2023). Hence, all the information about a class can be unlearned by forgetting only a subset of the forget class training data in the case of Unlearning via Examples, making it very data-efficient.
5.3 Comparison with Baseline
We now compare the results of both the proposed methods, which require Backbone + DKVB models against the SCRUB method, which is optimized for models without such a bottleneck. For this, we will use the Backbone + Linear Layer model described in 5.1. On this model, we run SCRUB and compare the performance drop after unlearning with the two proposed methods. [^2] We refer to Appendix E.2 for more details on the implementation of this baseline Table 2 shows the comparison between the two previously reported methods and the SCRUB baseline. We can see that the proposed methods perform at least as well as SCRUB; in the case of CIFAR-100 and LACUNA-100 favorably, as we see from the damage incurred on the retain set test performance for CIFAR-100 and LACUNA-100. Finally, it is important to re-emphasize that the proposed methods achieve the shown performance without requiring any additional gradient-based training for unlearning. SCRUB is stopped when the forget set is completely unlearned without damaging the performance on the retain set. We refer to Appendix E.3 for further training details.
Table 2: Comparison between the proposed methods and the baseline on CIFAR-10, CIFAR-100 and LACUNA-100. We compare the change in performance on the retain and forget test data relative to the originally trained models. All three methods are able to unlearn the forget sets completely in all three cases and maintain the performance on the retain set of the data in the case of CIFAR-10. For CIFAR-100 and LACUNA-100, the proposed methods are able to maintain the performance on the retain set better than the baseline.
CIFAR-10 CIFAR-100 LACUNA-100
Rel. change from base model $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$ $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$ $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$
DKVB via Activations (sec 5.2) 0.36% -100% -0.20% -100% -0.17% -100%
DKVB via Examples (sec 5.2) 0.45% -100% -0.36% -100% -0.09% -100%
Linear Layer + SCRUB 1.62% -100% -0.91% -100% -1.10% -100%
5.4 Unlearning beyond the zero-shot setting
@@
Next, we investigate the effect of using additional compute to the proposed methods. As shown previously, the proposed methods perform competitively to SCRUB zero-shot. To the best of our knowledge, SCRUB is the most competitive and relevant unlearning approach. However, it has the inherent drawback of requiring compute for unlearning. Nevertheless, for a fair comparison, we additionally explore the implications of this additional compute for the proposed two methods. Specifically, we retrain the models after zero-shot unlearning on the training data of the retain set (i.e., $\mathcal{D}_{train}^{retain}$) for 10 epochs. For the baseline, we use the same experimental setting as in Section 5.3 and run it for 10 epochs, instead of stopping once the forget set has been completely unlearned. Figure 4 highlights the effect of retraining of the proposed methods compared to SCRUB across multiple epochs, for all three datasets.
@@
Retraining the unlearned models on the retain set does not affect their performance significantly. The performance of the baseline on the other hand increases after an initial drop in case of CIFAR-100 and LACUNA-100. The initial drop may be attributed to the damage to the retain set performance caused by the initial max-steps. The subsequent increase can be attributed to the fact that the SCRUB training objective also optimizes the task loss on the retain set. Thus, once the model unlearns the forget set, SCRUB shifts the model capacity towards better learning the retain set. For CIFAR-10 this results in the model performing better than the DKVB models on the retain set as the retain set test accuracy after unlearning is higher than the original model. However, the baseline is unable to recover its original performance for CIFAR-100 and LACUNA-100. We refer to Appendix B for a similar comparison on the forget set.
@@
6 Limitations and Future Work
The proposed methods inherit the limitations of the DKVB (Träuble et al., 2023). Two important limitations are 1.) the reliance of DKVB on pre-trained encoders which can extract meaningful shared representations and 2.) trade-offs in downstream performance due to the use of an information bottleneck. Extensions to the model may involve training sparse representations inducing discrete bottleneck end-to-end. Further, in our experiments, we consider the setting of class incremental learning where the forget set can be easily identified and isolated. However, this may not always be true for a given task and more complicated approaches might be needed to identify the data that needs to be removed from the model. Thus, future work may involve scaling the proposed methods to such tasks and developing methods to identify and isolate the forget set in such cases.
7 Conclusion
In this work, we proposed a new approach to unlearning that requires negligible additional computation in order to unlearn a subset of data. This approach is based on the use of a discrete architectural bottleneck which induces sparse representations. These sparse representations facilitate unlearning a subset of data from the model with minimal to no performance drop on the rest of the data. We focused on the setting of class unlearning and our experiments show that the proposed approach, while being compute efficient, performs competitively with or in some cases better than a state-of-the-art approach which requires additional compute to perform unlearning. We also studied a compute-matched condition in which we allowed additional retraining. We found that the proposed approach did not benefit from such retraining, indicative of the fact that knowledge is highly localized in the Key-Value Bottleneck. Consequently, excising the activated key-value pairs from the model is a highly effective means of unlearning the forget set without disrupting the retain set.
Appendix A Deciding the Forget Class
We assume that this class should be the most difficult one for the model to forget. Figures 5 - 5 show the number of mis-classifications per class on the test data, for all three datasets. For CIFAR-10, class #1 is the best-learned class with the lowest number of mis-classifications. Thus, we select class #1 as the forget class for the dataset. Similarly, for CIFAR-100 class 58 is the best-learned class and for LACUNA-100, class 48 is one of the best-learned classes with zero mis-classifications. Hence, we select classes #58 and #48 as the forget classes for CIFAR-100 and LACUNA 100 respectively. We use the same forget classes for experiments on the models with a linear layer in place of the DKVB (i.e., the baseline) as well.
Appendix B Performance on Forget Class during Re-training
In all three cases, the baseline completely unlearns the forget set quickly.
@@ Appendix B Performance on Forget Class during Re-training
Appendix C Detailed Analysis Zero-Shot Unlearning via Activations
C.1 Unlearning via Activations
CIFAR-10: For CIFAR-10, we first validate the unlearning with $N_{a}\in\{10000n:n\in[0,20]\}$ and record the performance on $\mathcal{D}_{test}^{retain}$ and $\mathcal{D}_{test}^{forget}$. Figure 2(c) shows the scatterplot of forget set test accuracy vs. retain set test accuracy for different values of $N_{a}$. We start off with $N_{a}=0$, i.e., no unlearning, to $N_{a}=300000$. The initial forget and retain set test accuracies for CIFAR-10 as shown in Table 1 are 93.02% and 96.50% respectively. As $N_{a}$ increases, the forget class test accuracy decreases, slowly at first and rapidly later. The model reaches random accuracy (i.e. 10% for CIFAR-10) on the forget class test data at $N_{a}=150000$. At this point the retain set test accuracy is 92.97%. The model unlearns the forget class completely between $N_{a}=170000$ (0.4%) and $N_{a}=180000$ (0%). The retain set test accuracy at $N_{a}=180000$ is 92.94%, which is almost similar to the initial accuracy of 92.61%. Further increasing $N_{a}$ till 200000 leads to an increase in retain set test accuracy to 93%.
CIFAR-100 We vary $N_{a}$ for CIFAR-100 as $N_{a}\in\{5000n:n\in[0,18]\}$. Figure 2(d) shows the scatterplot of retain set test accuracy vs. forget set test accuracy for these values. We start with retain set test accuracy of 78.24% and forget set test accuracy of 96% (see Table 1. On increasing $N_{a}$, the retain set test accuracy fluctuates in a small range around the original accuracy, reaching a maximum of 78.32% at $N_{a}=40000$ and $450000$. The forget set test accuracy drops slowly at first, dropping by 12% till $N_{a}=20000$. The model reaches random test accuracy (1% for CIFAR-100) on the forget set at $N_{a}=65000$ where the forget set test accuracy and retain set test accuracies are 1% and 78.14% respectively. It reaches 0% at $N_{a}=70000$. At this point, the retain set test accuracy is 78.08%. Thus the model incurs a 0.2% deterioration in its retain set test accuracy by the time it has completely unlearned the forget class.
LACUNA-100 We vary $N_{a}$ for LACUNA-100 as $N_{a}\in\{10000n:n\in[0,20]\}$. Figure 2(e) shows the scatterplot of retain set test accuracy vs forget set test accuracy for LACUNA-100. The initial retain set test accuracy and forget set test accuracy for LACUNA-100 are 90.28% and 100% respectively (see Table 1. As we increase $N_{a}$, the retain set test accuracy increases monotonically to 90.43% till $N_{a}=40000$ and then decreases before starting to fluctuate. Unlike CIFAR-10 and CIFAR-100, the decrease in forget set test accuracy is rapid from the beginning and slows down towards the end. The forget set test accuracy reaches a near random accuracy of 2% at $N_{a}=140000$ where the retain set test accuracy is 90.24%. The model fully unlearns the forget set (i.e., forget set test accuracy = 0%) at $N_{a}=150000$ at which point the accuracy of the model on the retain set test data is 90.13%. Thus there is only a 0.17% drop in performance as compared to the original model. Moreover as seen in Figure 2(e), the retain set test accuracy increases slightly as $N_{a}$ is increased further.
C.2 Unlearning via Examples
CIFAR-10 CIFAR-10 contains 5000 examples in the forget set training data. We unlearn randomly sampled $N_{e}$ examples out of these 5000 examples. We experiment with the following values of $N_{e}\in\{200n:n\in[0,20]\}$. Figure 3(a) plots the forget set test performance vs. retain set test performance of the model for different values of $N_{e}$. We start off with retain set test and forget set test accuracies of 92.61% and 96.50% respectively. Similarly to in unlearning via activations, the forget class test accuracy decreases with an increase in $N_{e}$, and the decrease is slow at first and then rapid. The retain set test accuracy, on the other hand, increases monotonically, although marginally. The model achieves random accuracy on the forget class between $N_{e}=2400$ and $N_{e}=2600$, where the accuracies are 18.04% and 6.68% respectively. The accuracy on retain set test data at these points are 92.95% and 92.98% respectively. The forget set test accuracy drops to 0% (i.e., complete unlearning) between $N_{e}=3000$ (0.08%) and $N_{e}=3200$ (0%). The retain set test accuracies at these two points are 92.99% and 93.02%. Further increasing $N_{e}$ does not affect the retain set test accuracy much.
CIFAR-100 CIFAR-100 contains 500 examples of the forget class in the training data. We vary $N_{e}$ for CIFAR-100 as $N_{e}\in\{10n:n\in[0,15]\}$. Figure 3(b) shows the plot of forget set test accuracy vs. retain set test accuracy. The forget set test accuracy drops rapidly and monotonically from $N_{e}=30$ to $N_{e}=110$ where its value is 5% and the retain set test accuracy is 77.99%. It reaches the near random accuracy on 1.6% in the forget set test data at $N_{e}=120$ where the retain set test accuracy is 78.01%. The model completely unlearns the forget set at $N_{e}=150$ where the retain set test accuracy is 77.96%. The retain set test accuracy first decreases (almost monotonically) till $N_{e}=90$. It increases thereafter slightly, before decreasing again towards the end.
LACUNA-100 LACUNA-100 has 400 forget class training examples. We unlearn a randomly selected subset of size $N_{e}$ from it. $N_{e}$ is varied as $N_{e}\in\{20n:n\in[0,15]\}$. Figure 3(c) shows the plot of retain set test accuracy to forget set test accuracy for different $N_{e}$. As in the case of unlearning via examples and unlike CIFAR-10 and CIFAR-100, the forget set test accuracy decreases rapidly with $N_{e}$ initially till $N_{e}=60$; and rather slowly after that. The retain set test accuracy increases slightly at first and then starts fluctuating. The model reaches a random performance of 1% on the forget set test data at $N_{e}=260$. The retain set test performance at this point is 90.25%. Finally, the model unlearns the forget class completely at $N_{e}=300$, where the retain set test accuracy is 90.2%. Thus, the retain set performance drops by 0.09% by the time the model has completely unlearned the forget class.
Appendix D Using the proposed Methods against Membership Inference Attacks
@@
Depending on the application, complete unlearning of the forget set may not always be the final goal of unlearning. For several use cases such as removing information about corrupted data from the model or removing harmful biases exhibited by the model, maximal error on the forget set is desirable. However, for applications such as Differential Privacy, it is more desirable to achieve a forget set error which is similar to that of a model trained from scratch only on the retain set. Otherwise, it makes the unlearned model susceptible to Membership Inference Attacks (MIA) (Shokri et al., 2017). Although we do not explore this setting in this work, the proposed method can also be used for applications where complete unlearning is not desirable. This can be done by following a procedure similar to SCRUB+R (Kurmanji et al., 2023), wherein instead of selecting a particular model checkpoint, one can select the model corresponding to specific values of $N_{a}$ or $N_{e}$ such that the error on the forget set test data is similar to the reference point as defined in Kurmanji et al. (2023).
@@
First,it is important to clarify that the proposed approach is not a-priori suited for selective unlearning, i.e. the setting where we want the model to forget specific examples or a small subset of examples instead of removing the information about an entire class. The KV bottleneck induces clusters of representation, where the members of a particular cluster correspond to the representations belonging to the same class. When we try to unlearn the representations corresponding to one particular example belonging to a particular class, the KV bottleneck routes the selection to other (key-)representations within the same cluster. Since these representations also contain information about the same class as the examples we intend to unlearn, the model would still predict the class to be unlearnt.
@@
Due to the same reason our approach is also not designed for working against traditional Membership Inference attacks. According to the basic attacks setup as explained in Kurmanji et al. (2023), the objective is to obtain a model that has unlearnt a small subset of specific examples (i.e. selective unlearning) such that the loss of the model on the unlearnt subset of examples should be indistinguishable from loss on examples that the model never saw during training.
@@
Nevertheless, we attempt to modify the above setup to that of ”Class Membership Inference Attacks (CMIA)”. In CMIA, the aim is to defend against an attacker whose aim is to determine whether a model that has undergone unlearning ever saw a particular class as a part of its training data. Thus, we want the model to unlearn a particular class such that the losses/performance of the model on the unlearnt class is indistinguishable from a held-out class that the model never saw during its training. We describe the experimental setup and results below.
@@
Experimental Setup We perform the experiment for CIFAR10. We divide the dataset into training data ($D_{Train}$), validation data ($D_{Val}$) and test data ($D_{Test}$). Training Data consists of 4000 examples per class; validation and test data consist of 1000 examples per class. We first trained a model on the first 9 classes of CIFAR10. Thus, class number 10 is the held-out class. Next, we unlearn class 1 from the model using the Unlearning via Activations approach introduced in the paper. We unlearn the model until the loss of the model on the validation sets of the forget class and the held-out class are similar. In our experiments, we find that we reach this point at approximately $N_{a}=240000$. The loss $l(x,y)$ in our case would be the cross-entropy loss.
@@
Next, we label the losses corresponding to the validation and test set of the forget class as 1 and those corresponding to the validation and test set of the held-out class as 0. We train a binary classifier on the validation losses of the forget and held-out sets and evaluate it on the test losses. We follow a similar setting for the baseline model, where we obtain the model suitable for MIA defense by using SCRUB+R (Kurmanji et al., 2023). For a successful defense, we would want the accuracy of the classifier to be close to 50%, indicating that it is unable to distinguish between the unlearnt class and the held-out class. Same as Kurmanji et al. (2023), we use sklearn.logistic_regression as our attacker (the binary classifier). We call the approach described above Partial UvA (Partial Unlearning via Activations). We run experiments for 3 random seeds, and the mean of the attacker performance is reported. Note that a similar procedure can also be followed using Unlearning via Examples.
@@
Observations and Results: We report the results of the experiment described above in the table given below. We observe that although the baseline performs better, the proposed approach performs competitively, even though we have not intended to develop the method for this scenario.
@@
Table 3: Comparison on Class Membership Inference Attacks between the proposed approach and the baseline. A binary classifier is trained on the validation losses of the forget and held-out sets and is evaluated on the test losses. The proposed approach performs competitively to SCRUB + R/
@@
Appendix E Training Details and Hyperparameters
We do not use any data augmentation in any experiment. The transforms used are the same as CLIP (Radford et al., 2021) pretrained ViT/B-32 transforms.
E.1 Training Details and Hyperparameters for training the original DKVB Models
Table 4 shows the hyperparameters used for training the base DKVB models

### v2

@@ A.1 Initial Performances of the Models
A.2 Deciding the Forget Class
We assume that this class should be the most difficult one for the model to forget. Figures 4 - 4 show the number of mis-classifications per class on the test data, for CIFAR-10, CIFAR-100 and LACUNA-100 for the ViT-B/32 backbone. For CIFAR-10, class #1 is the best-learned class with the lowest number of mis-classifications. Thus, we select class #1 as the forget class for the dataset. For CIFAR-100 class 58 is the best-learned class and for LACUNA-100, class 48 is one of the best-learned classes with zero mis-classifications. Hence, we select classes #58 and #48 as the forget classes for CIFAR-100 and LACUNA 100 respectively. We determine the forget class in all other cases using the same method. We use the same forget classes for experiments on the models with a linear layer in place of the DKVB (i.e., the baseline) as well. Table 4 shows the forget class for all the cases discussed in our experiments.
Table 4: Forget classes for the different scenarios presented in the paper
Forget Classes CIFAR-10 CIFAR-100 LACUNA-100 ImageNet-1k
ViT-B/32 1 58 48 1
ResNet-50 1 94 34 9
A.3 Multi Class Unlearning
We attempt to investigate the effects of unlearning multiple classes at once by performing experiments on CIFAR-100 for both ViT/B-32 and ResNet-50 models. We unlearn upto 10 classes using both - Unlearning via Activations as well as Unlearning via Examples, and run each experiment for 5 seeds. The classes to be forgotten are chosen randomly. The given below show the unlearnt model’s performance on the retain class vs the number of classes unlearnt at approximately the point of complete unlearning. Figure 5 shows the results of this experiment.
Figure 5: Multi Class Unlearning for CIFAR-10
We can see that a ResNet-50 backbone performs comparatively better. Also, Unlearning via Activations performs relatively better as compared to Unlearning via Examples. We can also clearly see that the relative error as compared to the original model on the retain set performance are higher than single class unlearning, as expected. Noticeably, the performance starts to significantly deteriorate when forgetting ¿ 6 classes in the case of Unlearning via Examples, and ¿ 8 examples in the case of Unlearning via Activations for both the backbones.
A.4 Choosing the forget class randomly
To ensure that the effectiveness of the approach is not class specific, we perform experiments for CIFAR-100, where the class to be forgotten is randomly chosen, and compare the performance of the proposed approaches against SCRUB [2]. We run each experiment for 5 random seeds, wherein the forget class is randomly chosen for each seed. Rest of the experimental setup remains the same as described in Section 5 of the paper. We report the results in Table 5
Table 5: Selecting the forget class randomly for CIFAR-100
CIFAR-100
Backbone Method $\mathcal{D}_{test}^{retain}$ $\mathcal{D}_{test}^{forget}$
ViT/B-32 DKVB via Activations (sec 5.2) -0.27 $\pm$ 0.07 % -100% $\pm$ 0 %
DKVB via Examples (sec 5.2) -0.47 $\pm$ 0.03 % -100 $\pm$ 0 %
ResNet-50 DKVB via Activations (sec 5.2) 0.58 $\pm$ 0.12 % -100 $\pm$ 0 %
DKVB via Examples (sec 5.2) 0.28 $\pm$ 0.11 % -100 $\pm$ 0 %
Clearly, even with the forget class chosen randomly, the proposed approaches perform equally well as in the scenario where the forget class is fixed. This proves that the effectiveness of the proposed approaches is not class dependent.
A.5 Unlearning beyond the compute free setting
@@
We investigate the effect of using additional compute to the proposed methods. As shown previously, the proposed methods perform competitively to SCRUB. To the best of our knowledge, SCRUB is the most competitive and relevant unlearning approach. However, it has the inherent drawback of requiring compute for unlearning. Nevertheless, for a fair comparison, we additionally explore the implications of this additional compute for the proposed two methods for a ViT/B-32 backbone on CIFAR-10, CIFAR-100 and LACUNA-100. Specifically, we retrain the DKVB models after the (compute efficient) unlearning, on the training data of the retain set (i.e., $\mathcal{D}_{train}^{retain}$) for 10 epochs. For the baseline, we use the same experimental setting as in Section 5.2.1, except - we run it for 10 epochs instead of stopping when either the forget set has been completely unlearned or the performance has converged. Figure 6 and figure 7 highlight the effect of retraining of the proposed methods compared to SCRUB across multiple epochs, for all three datasets.
@@
Retraining the unlearned models on the retain set does not affect their performance significantly. The performance of the baseline on the other hand increases after an initial drop in case of CIFAR-100 and LACUNA-100. The initial drop may be attributed to the damage to the retain set performance caused by the initial max-steps. The subsequent increase can be attributed to the fact that the SCRUB training objective also optimizes the task loss on the retain set. Thus, once the model unlearns the forget set, SCRUB shifts the model capacity towards better learning the retain set. For CIFAR-10 this results in the model performing better than the DKVB models on the retain set as the retain set test accuracy after unlearning is higher than the original model. However, the baseline is unable to recover its original performance for CIFAR-100 and LACUNA-100.
@@
For the forget set, in all three cases, the baseline completely unlearns the forget set quickly within the first few epochs, as shown in Figure 7.
@@
A.6 Using the proposed Methods against Membership Inference Attacks
@@
Depending on the application, complete unlearning of the forget set may not always be the final goal of unlearning. For several use cases such as removing information about corrupted data from the model or removing harmful biases exhibited by the model, maximal error on the forget set is desirable. However, for applications such as Differential Privacy, it is more desirable to achieve a forget set error which is similar to that of a model trained from scratch only on the retain set. Otherwise, it makes the unlearned model susceptible to Membership Inference Attacks (MIA) (Shokri et al., 2017). Although we do not explore this setting in detail in this work, the proposed method can also be used for applications where complete unlearning is not desirable. This can be done by following a procedure similar to SCRUB+R (Kurmanji et al., 2023), wherein instead of selecting a particular model checkpoint, one can select the model corresponding to particular values of $N_{a}$ or $N_{e}$ such that the error on the forget set test data is similar to the reference point as defined in Kurmanji et al. (2023).
@@
First,it is important to clarify that the proposed approaches are not a-priori suited for selective unlearning, i.e. the setting where we want the model to forget specific examples or a small subset of examples instead of removing the information about an entire class. The KV bottleneck essentially induces induces clusters of representation, where the members of a particular cluster correspond to the representations belonging to the same class (see Figure 2 in Träuble et al. (2023)). When we try to unlearn the representations corresponding to one particular example belonging to a particular class, the KV bottleneck routes the selection to other (key-)representations within the same cluster since those keys would be the next closest to the representation of the encoder. Since these representations also contain information about the same class as the examples we intend to unlearn, the model would still predict the class to be unlearnt.
@@
Due to the same reason the proposed approaches are also not designed for working against traditional Membership Inference attacks. According to the basic attacks setup as explained in Kurmanji et al. (2023), the objective is to obtain a model that has unlearnt a small subset of specific examples (i.e. selective unlearning) such that the loss of the model on the unlearnt subset of examples should be indistinguishable from loss on examples that the model never saw during training.
@@
Nevertheless, since the proposed approaches are designed for class unlearning specifically, we attempt to evaluate them on a modified version of the above. We call this ”Class Membership Inference Attacks (CMIA)”. In CMIA, the aim is to defend against an attacker whose aim is to determine whether a model that has undergone unlearning ever saw a particular class as a part of its training data. Thus, we want the model to unlearn a particular class such that the losses/performance of the model on the unlearnt class as a whole, is indistinguishable from those on a held-out class that the model never saw during its training. We describe the experimental setup and results below.
@@
Experimental Setup We perform the experiment for CIFAR10 with a ViT/B-32 backbone. We divide the dataset into training data ($D_{Train}$), validation data ($D_{Val}$) and test data ($D_{Test}$). Training Data consists of 4000 examples per class; validation and test data consist of 1000 examples per class. We first trained a model on the first 9 classes of CIFAR10. Thus, class number 10 is the held-out class. Next, we unlearn class 1 from the model using the Unlearning via Activations approach introduced in the paper. We unlearn the model until the loss of the model on the validation sets of the forget class and the held-out class are similar. In our experiments, we find that we reach this point at approximately $N_{a}=240000$. The loss $l(x,y)$ in our case is be the cross-entropy loss.
@@
Next, we label the losses corresponding to the validation and test set of the forget class as 1 and those corresponding to the validation and test set of the held-out class as 0. We train a binary classifier on the validation losses of the forget and held-out sets and evaluate it on the test losses. We follow a similar setting for the baseline model, where we obtain the model suitable for MIA defense by using SCRUB+R (Kurmanji et al., 2023). For a successful defense, we would want the accuracy of the classifier to be close to 50% on the test losses, indicating that it is unable to distinguish between the unlearned class and the held-out class. Same as Kurmanji et al. (2023), we use sklearn.logistic_regression as our attacker (the binary classifier). We call the approach described above Partial UvA (Partial Unlearning via Activations). We run experiments for 3 random seeds, and the mean of the attacker performance is reported. Note that a similar procedure can also be followed using Unlearning via Examples.
@@
Observations and Results: We report the results of the experiment described above in the table given below. We observe that although the baseline performs slightly better, the proposed approaches perform competitively, even though we have not intended to develop the method for this scenario.
@@
Table 6: Comparison on Class Membership Inference Attacks between the proposed approach and the baseline. A binary classifier is trained on the validation losses of the forget and held-out sets and is evaluated on the test losses. The proposed approach performs competitively to SCRUB + R
@@
A.7 Training Details and Hyperparameters
We perform all of our experiments on a 48GB RTX8000 GPU. We do not use any data augmentation in any experiment. The transforms used for training the model with a ViT/B-32 backbone are the same as CLIP (Radford et al., 2021) pretrained ViT/B-32 transforms. For ResNet-50, both pre-trained weights and transforms are loaded from torchvision.models.ResNet50_Weights
A.7.1 Training Details and Hyperparameters for training the original DKVB Models
In the case of ImageNet pretrained ResNet-50, the representations of the backbone are extract from the 3rd last layer for CIFAR-10, CIFAR-100 and LACUNA-100 and from the 4th last layer for ImageNet-1k. Table 7 shows all the hyperparameters used for training the base DKVB models.